# Collaborative Perception for Connected and Autonomous Driving: Challenges, Possible Solutions and Opportunities

### v1

@@ Abstract
Autonomous driving has attracted significant attention from both academia and industries, which is expected to offer a safer and more efficient driving system. However, current autonomous driving systems are mostly based on a single vehicle, which has significant limitations which still poses threats to driving safety. Collaborative perception with connected and autonomous vehicles (CAVs) shows a promising solution to overcoming these limitations. In this article, we first identify the challenges of collaborative perception, such as data sharing asynchrony, data volume, and pose errors. Then, we discuss the possible solutions to address these challenges with various technologies, where the research opportunities are also elaborated. Furthermore, we propose a scheme to deal with communication efficiency and latency problems, which is a channel-aware collaborative perception framework to dynamically adjust the communication graph and minimize latency, thereby improving perception performance while increasing communication efficiency. Finally, we conduct experiments to demonstrate the effectiveness of our proposed scheme.
I Introduction
@@
With the development of artificial intelligence and computationally efficient hardware, autonomous driving has attracted significant attention from both academia and industries, which is expected to bring a safer and more efficient driving system [1]. However, current autonomous driving system is mostly based on a single vehicle, which has significant limitations that may threaten the driving safety. For example, occlusion is an inevitable problem that a single-agent perception system cannot address, which may cause serious accidents. In addition, the sensing range of a single-agent perception system is limited by the sensors (e.g., the maximum detection distance of LiDAR, the field of view of camera, etc.), which may miss some important information. As a result, we need to develop a more robust perception system to address these limitations. Recently, vehicle-to-vehicle (V2V) and vehicle-to-everything (V2X) communications have emerged as viable technologies for autonomous driving, which opens up a new field of research, namely, collaborative perception with connected and autonomous vehicles (CAVs). Collaborative perception with CAVs shows a promising solution to tackling the limitations on a single-vehicle perception system, which can improve the reliability and safety of autonomous driving. In addition, it presents a wealth of opportunities, including an improved user experience, heightened road safety, better air quality, and a range of innovative transportation solutions. However, to effectively implement this technology, several critical design issues need to be addressed. These include concerns about data privacy and ethics, the development of smart city infrastructure, standardization and regulatory policies from governments and regulatory bodies, building trust among humans and the system, and finding the right balance between over-reliance and communication overhead. Therefore, while the benefits of collaborative perception are abundant, a comprehensive and multifaceted approach is essential for its successful integration into everyday life.
@@
In order to investigate collaborative perception, some public large-scale datasets for collaborative perception have been released. For example, Xu et al. proposed OPV2V [2], which is an open dataset for collaborative perception, supporting LiDAR-based 3D object detection and camera-based bird’s eye view (BEV) segmentation. Li et al. proposed V2X-Sim [3], which is the first synthetic V2X-aided collaborative perception dataset in autonomous driving that supports object detection, tracking, and BEV semantic segmentation. Moreover, there are many studies dedicated to designing collaboration modules that aim to balance spectrum bandwidth and perception accuracy. These focus on addressing issues such as with whom to communicate, when to communicate, how to share information, and how to fuse and aggregate features, among others. However, these works assume an ideal scenario, which do not consider practical autonomous driving environments. In fact, in order to realize efficient, reliable, and safe collaborative perception, there are still many open challenges to be addressed, such as information sharing asynchrony, pose errors, and the collaboration security and privacy. In this article, we investigate the challenges of realizing collaborative perception and elaborate possible solutions based on various technologies, and we also point out some future research opportunities. Finally, we leverage an exemplar scheme to address some of the challenges in collaborative perception and conduct experiments to demonstrate the effectiveness of our proposed scheme.
This article is organized as follows. Sec. II provides a background of collaborative perception. In Sec. III, we discuss the challenges of realizing collaborative perception. In Sec. IV, we investigate the possible solutions based on various technologies and research opportunities. In Sec. V, we propose a scheme as an example to address the communication efficiency and transmission latency. Finally, in Sec. VI, we conduct experiments to demonstrate the effectiveness of our proposed scheme.
@@ II-A Demands for Autonomous Driving
A major factor driving the development of autonomous vehicles is the belief that they will reduce the number of accidents, injuries, and fatalities compared to vehicles driven by humans. However, many of the comparisons made so far lack fairness. While self-driving cars do not suffer from fatigue or emotional impairments like angers or frustrations, they do lack the ability to navigate uncertain and ambiguous situations with the same level of skill and foresight as an experienced and alert human driver.
Likewise, self-driving vehicles may make incorrect decisions because of errors in detecting and recognizing objects. This can be compared to the poor choices a human driver might make under the influence of alcohol. Decisions made by autonomous vehicles due to these detection failures could be as detrimental, if not more so, as those made by their human counterparts. For example, in a crash involving a Tesla in California, the vehicle made a critical error, leading to a fatal outcome: its sensors detected the concrete barrier but dismissed this information, mistakenly interpreting the barrier’s stationary position on the radar as non-threatening. Additionally, various errors made by autonomous vehicles have also led to accidents, resulting in loss of life and property. As a result, the safety of autonomous vehicles has become a significant public concern, underscoring the need for the development of safer autonomous driving.
@@ II-C Development of V2X Communications
The initial technologies for vehicular communications primarily encompasses vehicle-to-vehicle (V2V) and vehicle-to-infrastructure (V2I) modes. V2V communications enable vehicles to exchange information amongst themselves, whereas V2I communications facilitate information access via roadside units (RSUs). With the advent of Internet of Things (IoT) technologies, the concept of vehicle-to-everything (V2X) communications have emerged, which is crucial for enabling collaborative perception in autonomous driving. The V2X communications could be enabled by various technologies, such as edge computing-enabled V2X communications and the 6G-enabled V2X communications.
In modern collaborative perception systems, it is essential to store and process data collected from various sensors. While cloud computing and data centers, connected through V2X communications, are commonly used for this purpose, the latency is often too long for real-time applications. Edge computing, which processes data closer to the data source at the network’s edge, offers a solution by reducing latency and enhancing response time. Furthermore, 6G is set to play a pivotal role in future CAVs. 6G-enabled V2X communication promises ultra-low latency, high data transmission rates, and reliable connections, with greater energy and spectrum efficiency than previous generations of wireless communications.
@@ III Design Challenges
In order to develop a reliable collaborative perception system for autonomous driving, there are a lot of challenges that need to be addressed, such as large data volume, asynchronous information sharing, the collaboration security, data volume that need to be transmitted is very large, and the pose errors. In this section, we will discuss these challenges in detail.
@@
Large Data Volume. In autonomous vehicles equipped with different sensors, such as camera, LiDAR, and radar. the volume of sensing results (e.g., point clouds, image sequences, etc.) is very large. For example, according to KITTI dataset [4], each frame produced by 3D laser scanners contains around 100,000 points. The smallest scene captured in this dataset includes 114 frames, resulting in a total of more than 10 million points. In order to realize real-time, efficient collaboration, how to share the useful information in the large volume of data is a challenge to be addressed.
@@
Asynchronous Information Sharing. In an LTE-V2X communication system, the average latency can reach as much as 131.30 ms. Additionally, the differing latency times from nearby CAVs to ego CAV across various communication channels can cause serious problems with data fusion. Specifically, in a collaborative perception system, CAVs need to share information with ego CAV, due to the large volume of data and the limited transmission rate, there is a time-lapse from the moment that other CAVs start sending information to the moment that the ego CAV receives the information. This time-lapse results in information sharing in different time instants, causing the ego CAV to fuse information from different moments rather than at the same time instant. Through experimentation, it has been observed that this information sharing latency problem greatly impairs the performance of a collaborative perception system, resulting in performance even inferior to that of a single-agent perception system.
@@
Pose Errors. For effective collaboration, it is essential that multiple CAVs share accurate poses with each other. This necessitates the synchronization of their individual data within a uniform spatial coordinate system, forming the foundation for their cooperative efforts. However, in practical scenarios, the 6 Degrees of Freedom (DoF) pose estimated by each CAV’s localization system is not always precise. This lack of precision causes unintended errors in relative pose estimations, which can significantly diminish the collaborative effort.
@@
Collaboration Security. Collaboration security is another challenge that needs to be addressed. Compared with a single-agent perception system, a collaborative perception system faces higher susceptibility to adversarial attacks. In collaborative perception, each CAV uses an identical model to independently transform local sensor inputs into feature maps. These CAVs, part of the same collaborative network, then exchange these feature maps. Each CAV combines the received feature maps with its own, significantly enhancing performance in subsequent tasks. However, this exchange of feature maps between CAVs also opens up possibilities for adversaries to compromise the entire system. For instance, an attacker could engage in a man-in-the-middle attack, altering the feature maps sent to the target CAV, or a malevolent CAV might directly transmit tampered feature maps to a victim. Additionally, since humans cannot visually interpret the encoded feature maps, even minor modifications to these maps can go undetected, making such attacks particularly covert. This vulnerability hinders the deployment of advanced collaborative perception tools in safety-critical environments. For instance, in autonomous driving, any failure in perception can lead to severe outcomes, including property damage or, in the worst case, loss of human lives.
@@ IV Possible Solutions and Opportunities
In this section, we will discuss the possible solutions to address the aforementioned challenges, and we also investigate the future research opportunities, including multi-modal collaboration, generalization to real world scenarios and collaboration security issues.
@@ IV-A Possible Solutions
Early data fusion is the most straightforward method, but it has the highest communication overhead. The late data fusion has the lowest communication overhead, but it may result in information loss. Recent research [5] indicates that intermediate fusion may yield the optimal trade-off between perception accuracy and bandwidth consumption.
@@
There already exist some research along this line. Wang et al. [6] leveraged a fully-connected graph neural network (GNN) to aggregate the intermediate features received from other CAVs. They first calculate the relative poses between other CAVs and ego vehicle, then warp the intermediate features to the ego vehicle’s coordinate system and send them to the ego vehicle. Secondly, they use a mask-aware accumulation operation to aggregate the intermediate features while ensuring only overlapping fields-of-view is considered. Hu et al. [7] proposed a spatial confidence-aware message fusion, which leverages multi-head attention to fuse the corresponding features from multiple CAVs at each individual spatial location. A key aspect of this design is the use of spatial confidence maps. These maps show how confident each CAV is about its data at different locations. By including these feature maps, the system can better learn and focus on the most reliable information from each CAV, enhancing the overall decision-making process.
@@
Asynchronous Information Sharing. Due to the differences of sensing instants from different CAVs and the channel variations between collaborative CAVs and the ego vehicle, shared information may arrive at the ego vehicle asynchronously. This is a critical issue in collaborative perception, which may cause the loss of some information. The main idea to tackle this problem is to compensate the time delay for either sensing or communication. For example, Wang et al. [6] proposed a method to compensate the time delay. Specifically, they constructed a graph in which each node represents the state representation of a CAV. They first initialize the state of each node in this graph. For every node, the intermediate features received from other CAVs, the relative 6 DoF poses, and the time delay relative to the sensor time of the receiving vehicle are fed into a convolutional neural network (CNN). This process yields a time-delay-compensated representation, effectively adjusting for the discrepancies in data acquisition timing among CAVs. Lei et al. [8] proposed a latency-aware framework, which utilizes past collaboration data to concurrently estimate the present feature and the associated collaboration attention, both of which are unknown because various kinds of the latencies. This framework can dynamically adjust the asynchronous perceptual features from multiple agents to align with the same time stamp, thereby enhancing the robustness and effectiveness of collaborative perception.
@@
Pose Errors. In order to address pose errors, there are two main methods: 1) deep learning based methods, 2) pose correction methods. Xu et al. [9] proposed an attention mechanism to handle pose errors, named multiscale window attention. This approach employs a hierarchical structure of windows, each with a unique attention span. By incorporating windows of varying sizes, the robustness of detecting objects is significantly enhanced, especially in face of localization inaccuracies. The larger windows in this system are designed to gather extensive visual information, aiding in counteracting major localization errors. In contrast, the smaller windows focus on finer details, maintaining the integrity of local visual contexts. Subsequently, a split-attention mechanism is utilized to efficiently merge inputs from these diverse windows. This integration enables multiscale window attention to effectively manage various types of pose-related errors. Lu et al. [10] proposed a pose correction method before transmitting the information. Specifically, they constructed an agent-object pose graph and optimized it, which could align the relative pose relationships between CAVs and the detected objects in the scene, thereby promoting the pose consistency. In addition, they proposed a multiscale intermediate fusion strategy to aggregate the information at multiple spatial scales, which could further alleviate the impact of pose errors.
@@
Collaboration Security. Zhao et al. [11] proposed a malicious CAV detection method, which can detect the evasion attacks against object detectors that aim to degrade the performance of a machine learning system at inference time. In particular, this approach utilizes the dynamics of CAV collaboration to effectively identify and eliminate malicious CAVs within an ego CAV’s collaboration network. It employs a dual-testing strategy to regulate the false positive rate. The testing involves two separate hypothesis evaluations, focusing on the consistency between the ego CAV and the CAV under inspection, with the baseline assumption that they are “consistent”. They introduced two innovative detection metrics for these tests: 1) a match loss statistic assessing the alignment between the bounding box proposals from both CAVs, and 2) a collaborative reconstruction loss statistic evaluating the congruence of their fused feature maps. A CAV is deemed malicious if it is found “inconsistent” with the ego CAV based on these tests. This method is distinct from proactive strategies as it not only shields CAVs from potential adversarial threats but also aids in identifying the sources of such attacks, which has significant social implications. Methodologically, this method combines the strengths of both detection metrics while maintaining a controlled false positive rate, and it is designed to be expandable to incorporate additional detection statistics from future research.

### v2

@@ Abstract
Autonomous driving has attracted significant attention from both academia and industries, which is expected to offer a safer and more efficient driving system. However, current autonomous driving systems are mostly based on a single-agent perception, which has significant limitations, causing serious safety concerns. Collaborative perception with connected and autonomous vehicles (CAV) shows a promising solution to overcoming these limitations. In this article, we first identify the challenges of collaborative perception, such as data sharing asynchrony, large data volume, and pose errors. Then, we discuss the possible solutions to address these challenges with various technologies, where the research opportunities are also elaborated. Furthermore, we propose a scheme to deal with communication efficiency and latency problems, which is a channel-aware collaborative perception framework to dynamically adjust the communication graph and minimize latency, thereby improving perception performance while increasing communication efficiency. Finally, we conduct experiments to demonstrate the effectiveness of our proposed scheme.
Index Terms:
Collaborative perception, autonomous driving, connected and autonomous vehicle (CAV), vehicle-to-everything (V2X) communication.I Introduction
@@
With the development of artificial intelligence and computationally efficient hardware, autonomous driving has attracted significant attention from both academia and industries, which is expected to bring a safer and more efficient driving system hanCollaborativePerceptionAutonomous2023a ; fang2024pacp . However, current autonomous driving system is mostly based on a single vehicle, i.e., a single-agent perception system, which has significant limitations that may threaten the driving safety. For example, occlusion is an inevitable problem that a single-agent perception system cannot overcome, which may cause serious accidents. In addition, the sensing range of a single-agent perception system is limited by the deployed sensors (e.g., the maximum detection distance of LiDAR and the field of view of camera), which may miss some important information. As a result, we need to develop a more robust perception system to address these limitations. Recently, vehicle-to-vehicle (V2V) and vehicle-to-everything (V2X) communications have emerged as viable technologies for autonomous driving, which opens up a new field of research, namely, collaborative perception with connected and autonomous vehicles. Collaborative perception with CAVs shows a promising solution to tackling the limitations on a single-agent perception system, which can improve the reliability and safety of autonomous driving. In addition, it presents a wealth of opportunities, including an improved user experience, heightened road safety, better air quality, and a range of innovative transportation solutions. However, to effectively implement this technology, several critical design issues need to be addressed. These include concerns about data privacy and ethics, the development of smart city infrastructure, standardization and regulatory policies from governments and regulatory bodies, building trust among humans and the system, and finding the right balance between over-reliance and communication overhead. Therefore, while the benefits of collaborative perception are abundant, a comprehensive and multifaceted approach is essential for its successful integration into everyday life.
@@
In order to investigate collaborative perception, some public large-scale datasets for collaborative perception have been released. For example, Xu et al. proposed OPV2V xuOPV2VOpenBenchmark2022 , which is an open dataset for collaborative perception, supporting LiDAR-based 3D object detection and camera-based bird’s eye view (BEV) segmentation. Moreover, there are many studies dedicated to designing collaboration modules that aim to balance spectrum bandwidth and perception accuracy. These focus on addressing issues such as with whom to communicate, when to communicate, how to share information, and how to fuse and aggregate features, among others. However, these works typically assume an ideal scenario that does not consider practical autonomous driving environments. For example, Where2comm huWhere2commCommunicationefficientCollaborative2024 assumes that information sharing is synchronous, but in practice, it may be asynchronous. In addition, most of these works assume that all the vehicles are benign huWhere2commCommunicationefficientCollaborative2024 ; wangV2VNetVehicletoVehicleCommunication2020a ; xuV2XViTVehicletoEverythingCooperative2022a , but in practice, some vehicles may be compromised, then become malicious and attack the collaborative perception system. These practical issues pose serious challenges for realizing reliable collaborative perception. In this article, we investigate the challenges in collaborative perception and elaborate possible solutions based on various technologies, and we also point out some future research opportunities. Finally, we leverage an exemplar scheme to address some of the challenges in collaborative perception and conduct experiments to demonstrate the effectiveness of our proposed scheme.
@@ II-A Demands for Autonomous Driving
A major factor driving the development of autonomous vehicles is the belief that they will reduce the number of accidents, injuries, and fatalities compared to vehicles driven by humans. However, self-driving autonomous vehicles may make incorrect decisions because of errors in detecting and recognizing objects. This can be compared to the poor choices a human driver might make under the influence of alcohol or fatigue. Decisions made by autonomous vehicles due to these detection failures could be as detrimental, if not more so, as those made by their human counterparts. For example, in a crash involving a Tesla in California, the vehicle made a critical error, resulting in a fatal outcome: its sensors detected the concrete barrier but dismissed this information, mistakenly interpreting the barrier’s stationary position on the radar as non-threatening. Additionally, various errors made by autonomous vehicles have also led to accidents, resulting in loss of life and property. As a result, the safety of autonomous driving has become a significant public concern, underscoring the need for the development of safer autonomous driving.
@@ II-C Development of V2X Communications
The initial technologies for vehicular communications primarily encompass vehicle-to-vehicle (V2V) and vehicle-to-infrastructure (V2I) modes. V2V communications enable vehicles to exchange information amongst themselves, whereas V2I communications facilitate information access via roadside units (RSUs). With the advent of Internet of Things (IoT) technologies, the concept of vehicle-to-everything (V2X) communications has emerged, which is crucial for enabling collaborative perception in autonomous driving. For example, dedicated short-range communications (DSRC) technology is specifically designed for vehicular communications, which can enable V2V and V2I communications. In addition, 5G NR-V2X is another promising technology to support both cellular and direct communications, including sidelink communications. However, the development of V2X communications faces several challenges, such as high mobility, dynamic network topology, and frequent handovers, which can result in high latency and packet loss. In order to tackle these challenges, 6G is set to play a pivotal role in future CAVs. 6G-enabled V2X communication promises to provide services with ultra-low latency, high data transmission rates, and reliable connections, with greater energy and spectrum efficiency than previous generations of wireless communications.
@@ III Design Challenges
In order to develop a reliable collaborative perception system for autonomous driving, there are a lot of design challenges that need to be addressed, such as large data volume, asynchronous information sharing, collaborative security, and pose errors. In this section, we will discuss these challenges in detail.
@@
Large Data Volume. In autonomous vehicles equipped with different sensors, such as camera, LiDAR, and radar. the volume of sensing results (e.g., point clouds, image sequences, etc.) is very large. For example, according to KITTI dataset geigerAreWeReady2012 , each frame produced by 3D laser scanners contains around 100,000 points. The smallest scene captured in this dataset includes 114 frames, resulting in a total of more than 10 million points. In order to realize real-time, efficient collaboration, how to share the useful information in the large volume of data is a grand challenge.
@@
Asynchronous Information Sharing. In an LTE-V2X communication system, the average latency can reach as much as 131.30 ms. Additionally, the differing latency from nearby CAVs to ego CAV across various communication channels can cause serious problems with data fusion. Specifically, in a collaborative perception system, CAVs need to share information with ego CAV, due to the large volume of data and the limited transmission rate, there is a time-lapse from the moment that other CAVs start sending information to the moment that the ego CAV receives the information. This time-lapse results in information sharing at different time instants, causing the ego CAV to fuse information at different instants rather than at the same time. Through experimentation, it has been observed that this information sharing latency problem greatly impairs the performance of a collaborative perception system, resulting in performance even inferior to that of a single-agent perception system.
@@
Pose Errors. For effective collaboration, it is essential that multiple CAVs share accurate poses with each other. This necessitates the synchronization of their individual data within a uniform spatial coordinate system, forming the foundation for their cooperative efforts. However, in practical scenarios, the six Degrees of Freedom (DoF) pose estimated by each CAV’s localization system is not always precise. This lack of precision causes unintended errors in relative pose estimation, which can significantly diminish the collaborative effort.
@@
Collaboration Security. Collaboration security is another challenge that needs to be addressed. Compared with a single-agent perception system, a collaborative perception system faces higher susceptibility to adversarial attacks. In collaborative perception, each CAV uses an identical model to independently transform local sensor inputs into feature maps. These CAVs, part of the same collaborative network, then exchange these feature maps. Each CAV combines the received feature maps with its own, significantly enhancing performance in subsequent tasks. However, this exchange of feature maps between CAVs also opens up possibilities for adversaries to compromise the entire system. For instance, an attacker could engage in a man-in-the-middle attack, altering the feature maps sent to the target CAV, or a malevolent CAV might directly transmit tampered feature maps to a victim. This is particularly difficult to detect if the feature maps are generated from adversarial perceptual information such as high-definition maps. Additionally, since humans cannot visually interpret the encoded feature maps, even minor modifications to these maps can go undetected, making such attacks particularly covert. This vulnerability hinders the deployment of advanced collaborative perception tools in safety-critical environments. For instance, in autonomous driving, any failure in perception can lead to severe outcomes, including property damage or, in the worst case, loss of human lives.
@@ IV Possible Solutions and Opportunities
In this section, we will discuss the possible solutions to address the aforementioned challenges, and we also investigate the future research opportunities, including multi-modal collaboration, generalization to real world scenarios, and collaboration security issues.
@@ IV-A Possible Solutions
Early data fusion is the most straightforward method, but it has the highest communication overhead. The late data fusion has the lowest communication overhead, but it may result in information loss. Recent research huAdaptiveCommunicationsCollaborative2023 indicates that intermediate fusion may yield the optimal trade-off between perception accuracy and communications.
@@
There already exist some research along this line. Wang et al. wangV2VNetVehicletoVehicleCommunication2020a leveraged a fully-connected graph neural network (GNN) to aggregate the intermediate features received from other CAVs. They first calculate the relative poses between other CAVs and ego vehicle, then warp the intermediate features to the ego vehicle’s coordinate system and send them to the ego vehicle. Secondly, they use a mask-aware accumulation operation to aggregate the intermediate features while ensuring only overlapping fields-of-view is considered. Hu et al. huWhere2commCommunicationefficientCollaborative2024 proposed a spatial confidence-aware message fusion, which leverages multi-head attention to fuse the corresponding features from multiple CAVs at each individual spatial location. A key aspect of this design is the use of spatial confidence maps. These maps show how confident each CAV is about its data at different locations. By including these feature maps, the system can better learn and focus on the most reliable information from each CAV, enhancing the overall decision-making process.
@@
Asynchronous Information Sharing. Due to the differences of sensing instants from different CAVs and the channel variations between collaborative CAVs and the ego vehicle, shared information may arrive at the ego vehicle asynchronously. This is a critical issue in collaborative perception, which may cause the loss of some information. The main idea to tackle this problem is to compensate the time delay for either sensing or communication. For example, Wang et al. wangV2VNetVehicletoVehicleCommunication2020a proposed a method to compensate the time delay. Specifically, they constructed a graph in which each node represents the state representation of a CAV. They first initialize the state of each node in this graph. For every node, the intermediate features received from other CAVs, the relative six DoF poses, and the time delay relative to the sensing time of the receiving vehicle are fed into a convolutional neural network (CNN). This process yields a time-delay-compensated representation, effectively adjusting for the discrepancies in data acquisition timing among CAVs. Lei et al. leiLatencyAwareCollaborativePerception2022a proposed a latency-aware framework, which utilizes past collaboration data to concurrently estimate the present feature and the associated collaboration attention, both of which are unknown because various kinds of the latencies. This framework can dynamically adjust the asynchronous perceptual features from multiple agents to align with the same time stamp, thereby enhancing the robustness and effectiveness of collaborative perception.
@@
Pose Errors. In order to address pose errors, there are two main methods: 1) deep learning based methods, 2) pose correction methods. Xu et al. xuV2XViTVehicletoEverythingCooperative2022a proposed an attention mechanism to handle pose errors, named multiscale window attention. This approach employs a hierarchical structure of windows, each with a unique attention span. By incorporating windows of varying sizes, the robustness of detecting objects is significantly enhanced, especially in face of localization inaccuracies. Lu et al. luRobustCollaborative3D2023a proposed a pose correction method before transmitting the information. Specifically, they constructed an agent-object pose graph and optimized it, which could align the relative pose relationships between CAVs and the detected objects in the scene, thereby promoting the pose consistency. In addition, they proposed a multiscale intermediate fusion strategy to aggregate the information at multiple spatial scales, which could further alleviate the impact of pose errors.
@@
Collaboration Security. Zhao et al. zhaoMaliciousAgentDetection2023 proposed a malicious CAV detection method, which can detect the evasion attacks against object detectors that aim to degrade the performance of a machine learning system at inference time. In particular, this approach utilizes the dynamics of CAV collaboration to effectively identify and eliminate malicious CAVs within an ego CAV’s collaboration network. It employs a dual-testing strategy to regulate the false positive rate. The testing involves two separate hypothesis evaluations, focusing on the consistency between the ego CAV and the CAV under inspection, with the baseline assumption that they are “consistent”. They introduced two innovative detection metrics for these tests: 1) a match loss statistic assessing the alignment between the bounding box proposals from both CAVs, and 2) a collaborative reconstruction loss statistic evaluating the congruence of their fused feature maps. A CAV is deemed malicious if it is found “inconsistent” with the ego CAV based on these tests. This method is distinct from proactive strategies as it not only shields CAVs from potential adversarial threats but also aids in identifying the sources of such attacks, which has significant social implications. Methodologically, this method combines the strengths of both detection metrics while maintaining a controlled false positive rate, and it is designed to be expandable to incorporate additional detection statistics from future research.